OpenStudy (anonymous):

Harry can rake the leaves in the yard 8 hours faster than his little brother Jimmy can. If they work together, they can complete the job in 3 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jimmy to complete this job on his own.

hero (hero):

The working together formula: \[\frac{M \times N}{M + N} = t\] M = time it takes one person to complete a job N = time it takes another person to complete the same job t = time it takes both to complete the job while working together

hero (hero):

In this case let's represent the situation by the following: \[\frac{H \times J}{H + J} = 3\] H = hours it takes Harry to complete the job while working alone J = hours it takes Jimmy to complete the job while working alone 3 = hours it takes both Harry and Jimmy to complete the job while working together.

OpenStudy (anonymous):

@Hero wow thank you so much for the help, greatly appreciated!!!

hero (hero):

Harry can do the job 8 hours faster than Jimmy, so the situation can be represented by H = J - 8 So replacing H with J - 8 in the equation yields: \[\frac{(J - 8) \times J}{(J - 8) + J} = 3\] Perform the operations on the left side to get \[\frac{J^2 - 8J}{2J - 8} = 3\] Multiply both sides by 2J - 8 to get \[J^2 - 8J = 3(2J - 8)\] Distribute on the right side to get: \[J^2 - 8J = 6J - 24\] Move everything to the right side: \[J^2 - 8J - 6J + 24 = 0\] \[J^2 - 14J + 24 = 0\] Now you have a factorable quadratic that remains. Factor the quadratic and continue solving for J.